# Multimodal Representation Learning and Fusion thanks:

### v1

@@ X Evaluation Metrics and Benchmarks
Evaluating multi-modal models is challenging due to messy and diverse data [Liang et al., 2021]. Benchmarks like MultiBench and MM-BigBench test generalization and robustness [Liang et al., 2021, Yang et al., 2023]. Missing modalities require diagnostic and intervention methods [Wu et al., 2024], often using fusion strategies like (4). Standardized metrics and comprehensive benchmarks remain essential to ensure fair comparisons across tasks and domains.
That’s why recent studies have begun offering us some new tools. One highlight is MultiBench [Liang et al., 2021]. This is not just any benchmark, it is a large-scale, unified setup built to test representation learning across 15 datasets, 10 modalities, 20 prediction tasks, and 6 different application areas. Sounds intense? That’s the point. It helps us see how well a model can generalize, how efficient it is in terms of time and memory, and how robust it remains when one modality, say, vision or text, suddenly gets weak or disappears altogether. These tests do not just give scores, they reveal true strength.
Another landmark is MM-BigBench [Yang et al., 2023], which takes a wider angle. It builds a kind of testbed for checking how models handle multi-modal content-understanding tasks. And it does not just look at accuracy. It brings in a bunch of metrics: Best Performance, Mean Relative Gain, Stability, Adaptability, a full 360-degree view. The message is pretty clear: one single number is not enough anymore. We need metrics in bunches, because a multi-modal system is complex and multi-skilled. You miss the full picture if you just look at the top-line score. Adding to that, a review from [Li et al., 2024b] surveys 211 benchmarks covering MLLMs across four core abilities, understanding, reasoning, generation, and application. Again, the call is for a clear, shared rulebook. Without that, we are just throwing darts in the dark.
@@
Another hot topic? Missing modalities. This is not a rare problem, it happens all the time. Real-world data? Always broken in some way. Maybe a sensor fails. Maybe half the input is corrupted. Papers like [Wu et al., 2024][McKinzie et al., 2023] zoom in on this. They explore both diagnostic tools, how to detect when a modality is gone, and intervention methods, which are tricks to make the system stay strong anyway. These studies remind us: if your multi-modal model only works when all inputs are perfect, then it is not really ready for real life.
Now, if you step back and look across all this literature, a loud voice echoes: we need better benchmarks, and we need better metrics. Public datasets are crucial, not just for transparency, but for actual science. We want repeatable experiments. We want fair, apples-to-apples comparisons between models. That’s why works like [Shi et al., 2023][Li et al., 2024a] argue we must go deeper, not just check surface-level matching, but really dig into cross-modal understanding. Does the model just align text to image, or does it actually “get” the connection?
@@
From a technical point of view, most of the models are still deep nets, CNNs, RNNs, and especially Transformers. Researchers also explore the usual suspects: early fusion, late fusion, and middle fusion, all tested in works like [Liang et al., 2021][Yang et al., 2023]. When it comes to testing stability, some follow the robustness-checking protocols from [McKinzie et al., 2023], where they purposely inject noise or drop out parts of the modality data and watch how the models behave.
@@ XI Heterogeneous Data, Missing Modalities, and Adversarial Attacks
Heterogeneous data and missing modalities pose significant challenges in multi-modal learning [Wu et al., 2024]. Factorized representations focus on shared structures when data is incomplete [Tsai et al., 2019].
@@
where $\mathcal{M}_{\text{avail}}$ is the set of available modalities, $\mathbf{W}_{m}$ is the projection matrix, and $\mathbf{z}$ is the robust representation. Asymmetric reinforcement prioritizes stable modalities [Gao et al., 2025].
@@
in this equation, $\mathcal{L}_{m}$ is the loss for modality $m$, $w_{m}$ is its weight, and $\mathcal{L}_{\text{impute}}$ is the imputation loss. Adversarial attacks require robust defenses [Garcia et al., 2019].
@@
Over the past few years, multi-modal learning has made impressive progress. We’ve seen strong performance in a wide range of domains, computer vision, natural language processing, healthcare, and more [Wu et al., 2024][Tsai et al., 2019]. But still, let’s be honest: this field is far from “solved.” Several deep-rooted challenges continue to limit how far we can go. Some of them are technical; others are more about how the real world works. Either way, we need to face them head-on.
@@
One of the biggest headaches is dealing with heterogeneous data, that is, information coming from totally different sources, each with its own format and structure [Vishwamitra et al., 2021]. Imagine a dataset combining text, images, and audio. Each of these modalities has its own rhythm, its own kind of noise, its own “language.” And to expect a single model to handle them all smoothly? It is hard. The model must learn not just from each modality but also across them, and that’s where things start to break down.
@@
Another major challenge? Missing modalities. And trust me, this is not rare. A sensor might fail. Data gets corrupted. Or, in some cases, certain signals are deliberately excluded for privacy or cost reasons [Wu et al., 2024][Ma et al., 2021]. The result? A performance drop, sometimes a big one. So now, the question becomes: how do we build models that still function when part of the input goes missing? Some recent solutions offer promising ideas. For example, factorized representations help models focus on shared structures even when data is incomplete [Tsai et al., 2019]. Others propose asymmetric reinforcement [Gao et al., 2025], guiding the model to rely more on stable signals. And then there are lightweight adaptation techniques [Reza et al., 2023], which offer practical, less resource-intensive fixes. These methods show one thing clearly: resilience is now a must.
@@
And resilience is not just about noise or gaps, it is also about adversarial attacks [Garcia et al., 2019]. In high-stakes environments like medical AI or autonomous driving, even tiny, human-imperceptible changes to input data can lead to completely wrong predictions. That’s scary. So, defending against such attacks, detecting them, resisting them, recovering from them, has become a core concern [Xu et al., 2024]. Multi-modal models need to be not just smart, but safe and secure.
@@
Then there is the issue of modality imbalance, a subtle but serious problem. Sometimes one modality dominates the learning process because its signals are stronger or easier to model. This leads to the model “ignoring” weaker but still-important signals [Cheng et al., 2022][Vishwamitra et al., 2021]. And that’s bad, because it reduces the diversity of information the model actually uses. So, what do we do? Some researchers suggest greedy modality selection [Cheng et al., 2022], where the model picks the most useful signals step by step. Others turn to game-theoretic regularization [Kontras et al., 2024], where learning is balanced like a strategic game between modalities. Both methods aim to reduce bias and boost fairness in training, something increasingly critical in today’s AI landscape.
@@
Looking across all this work, one thing is obvious: solving these core issues is key to building stronger, more reliable multi-modal systems [Wu et al., 2024][McKinzie et al., 2023]. It is not just about improving one metric here or there. It is about improving how we learn joint representations, how we deal with data loss, and how we protect against manipulation. And interestingly, many of the proposed techniques are cross-cutting. For example, factorized learning and asymmetric reinforcement are useful not just for missing data, but also for handling imbalance and improving robustness overall [Tsai et al., 2019][Gao et al., 2025]. That kind of versatility is powerful.
@@
In summary, while the field of multi-modal learning has achieved notable progress, a range of persistent challenges continues to limit its broader adoption and effectiveness. Key technical obstacles, such as the integration of heterogeneous data, resilience to missing or corrupted modalities, and defense against adversarial interference, remain unsolved. Existing approaches, including factorized learning frameworks, robust training techniques, and lightweight adaptation methods, have demonstrated value across multiple problem domains rather than in isolation. This suggests that future research should focus on developing integrated solutions capable of addressing overlapping challenges within a unified framework. By understanding the interconnections between these problems, and by designing models that reflect such complexity, the field can move closer to producing truly robust, adaptive, and generalizable multi-modal systems suitable for deployment in real-world, imperfect environments [Wu et al., 2024][McKinzie et al., 2023].
@@ XII Future Directions in Multi-modal Research
As multi-modal learning continues to grow, clearer frameworks and practical taxonomies are increasingly critical [Cohn et al., 2024, Liang et al., 2024c]. These tools help researchers manage the complexity of heterogeneous data and ambiguous integration rules [Baltrušaitis et al., 2018, Rahate et al., 2022]. Structuring the field is as vital as advancing model design.
A key direction is the push for task-relative definitions of multi-modality, focusing on modalities relevant to specific tasks [Parcalabescu et al., 2021]. This approach enhances efficiency by prioritizing utility over exhaustive modeling, though challenges arise with unbalanced or misaligned modalities [Li and Tang, 2024, Munikoti et al., 2024]. Deep learning for co-learning and robust fusion strategies, such as (4) and (5), are gaining traction [Akkus et al., 2023], enabling modalities to guide each other in challenging settings [Wu et al., 2023].
Despite progress, challenges persist. Combining diverse data types is inherently complex, with issues like unstructured, noisy, or missing data [Wu et al., 2024]. Large-scale labeled datasets, particularly for non-text modalities, remain scarce [Liang et al., 2021, He et al., 2024]. These gaps hinder training flexible and reliable models, necessitating adaptive systems for imperfect conditions.
The field’s interdisciplinary nature, blending computer science, cognitive research, and beyond, will drive future growth. Flexible, real-world-ready solutions are essential, balancing theoretical advances with practical constraints like limited resources and incomplete labels. Ultimately, improving definitions, fusion strategies, and evaluation tools without overcomplicating systems will unlock multi-modal learning’s potential in AI, healthcare, and education [Cohn et al., 2024].
@@ XIII The Evolving Landscape of Multi-modal Learning
The field of multi-modal learning has gone through noticeable changes in recent years. This shift has been largely driven by the rapid progress in areas like machine learning, computer vision, and natural language processing. According to studies such as [Cohn et al., 2024][Parcalabescu et al., 2021][Baltrušaitis et al., 2018][Liang et al., 2024c][Akkus et al., 2023], the field has grown into a dynamic research hotspot, with a wide range of tasks and challenges emerging along the way. A shared belief among many researchers is that combining different types of data, text, image, audio, or beyond, can offer a more complete and nuanced understanding of complex problems [Rahate et al., 2022][Lee et al., 2023]. This idea has inspired the creation of new definitions, conceptual models, and frameworks to address the unique challenges that arise when working with multi-modal inputs [Parcalabescu et al., 2021][Lu, 2023].
@@
Across the literature, some recurring patterns begin to surface. A common one is the difficulty of handling multiple modalities: not just processing them in isolation, but figuring out how to connect them, align them, and model their interactions in a meaningful way [Baltrušaitis et al., 2018][Liang et al., 2024c]. At the same time, more and more researchers are turning to deep learning as a backbone for these efforts, applying it across domains such as healthcare, education, and autonomous systems [Akkus et al., 2023][McKinzie et al., 2023]. In parallel, identifying open questions and mapping future directions has become a key feature of many recent surveys and reviews [Liang et al., 2024c][Pattnayak et al., 2024], a sign that the field is still very much evolving.
@@
Interestingly, there is no complete consensus on what multi-modality even means. Some researchers propose their own definitions, some build entirely new taxonomies, and others focus on different fusion strategies, early, late, mid-level, or even graph-based approaches. For instance, [Cohn et al., 2024] provides a comprehensive literature review, while [Parcalabescu et al., 2021] introduces a redefined concept of modality interaction. These variations do not contradict each other so much as reflect the broad and multifaceted nature of the field.
@@
From a technical standpoint, the methods explored in these papers are also quite diverse. Some focus on theory and modeling, while others are more application-oriented. For example, [Abdollahzadeh et al., 2021] investigates meta-learning through a multi-task lens, and [Shi and Niethammer, 2023] proposes a strategy that alternates between increasing and decreasing inter-modal correlations to improve learning outcomes. All of this suggests that multi-modal learning is not a single discipline but a meeting point of many others, requiring interdisciplinary knowledge and flexibility in thinking.
@@
Still, despite all the enthusiasm, several core difficulties remain. There are calls for deeper theoretical work to solidify the foundations of this area [Liang et al., 2024c][Lu, 2023]. Mixing and managing different types of data continues to be a technical bottleneck [Baltrušaitis et al., 2018][Akkus et al., 2023]. Moreover, training large multi-modal models often demands significant computational resources and access to large-scale, high-quality datasets [Akkus et al., 2023][Pattnayak et al., 2024]. Another concern that comes up again and again is the lack of transparency: many current models are still hard to interpret or explain in practice [Liang et al., 2024c][Lee et al., 2023], which limits trust and usability in real-world settings.
@@
Overall, multi-modal learning has come a long way, and it still has a long way to go. Thanks to developments in machine learning and related fields, the tools we now have are more powerful and versatile than ever. But alongside this progress are real challenges, both technical and conceptual. What makes this area especially exciting is its potential: the promise of building smarter, more robust, and more interpretable models. With continued research and refinement, we can expect to see new breakthroughs and broader applications, especially in areas like computer vision, natural language processing, and human-computer interaction [Rahate et al., 2022][Lee et al., 2023][Cohn et al., 2024].

### v2

@@ X Evaluation Metrics and Benchmarks
Evaluating multi-modal models is challenging due to messy and diverse data [2107.07502]. Benchmarks like MultiBench and MM-BigBench test generalization and robustness [2107.07502, 2310.09036]. Missing modalities require diagnostic and intervention methods [2409.07825], often using fusion strategies like (4). Standardized metrics and comprehensive benchmarks remain essential to ensure fair comparisons across tasks and domains.
That’s why recent studies have begun offering us some new tools. One highlight is MultiBench [2107.07502]. This is not just any benchmark, it is a large-scale, unified setup built to test representation learning across 15 datasets, 10 modalities, 20 prediction tasks, and 6 different application areas. Sounds intense? That’s the point. It helps us see how well a model can generalize, how efficient it is in terms of time and memory, and how robust it remains when one modality, say, vision or text, suddenly gets weak or disappears altogether. These tests do not just give scores, they reveal true strength.
Another landmark is MM-BigBench [2310.09036], which takes a wider angle. It builds a kind of testbed for checking how models handle multi-modal content-understanding tasks. And it does not just look at accuracy. It brings in a bunch of metrics: Best Performance, Mean Relative Gain, Stability, Adaptability, a full 360-degree view. The message is pretty clear: one single number is not enough anymore. We need metrics in bunches, because a multi-modal system is complex and multi-skilled. You miss the full picture if you just look at the top-line score. Adding to that, a review from [2409.18142] surveys 211 benchmarks covering MLLMs across four core abilities, understanding, reasoning, generation, and application. Again, the call is for a clear, shared rulebook. Without that, we are just throwing darts in the dark.
@@
Another hot topic? Missing modalities. This is not a rare problem, it happens all the time. Real-world data? Always broken in some way. Maybe a sensor fails. Maybe half the input is corrupted. Papers like [2409.07825][2304.04385] zoom in on this. They explore both diagnostic tools, how to detect when a modality is gone, and intervention methods, which are tricks to make the system stay strong anyway. These studies remind us: if your multi-modal model only works when all inputs are perfect, then it is not really ready for real life.
Now, if you step back and look across all this literature, a loud voice echoes: we need better benchmarks, and we need better metrics. Public datasets are crucial, not just for transparency, but for actual science. We want repeatable experiments. We want fair, apples-to-apples comparisons between models. That’s why works like [2311.02692][2408.08632] argue we must go deeper, not just check surface-level matching, but really dig into cross-modal understanding. Does the model just align text to image, or does it actually “get” the connection?
@@
From a technical point of view, most of the models are still deep nets, CNNs, RNNs, and especially Transformers. Researchers also explore the usual suspects: early fusion, late fusion, and middle fusion, all tested in works like [2107.07502][2310.09036]. When it comes to testing stability, some follow the robustness-checking protocols from [2304.04385], where they purposely inject noise or drop out parts of the modality data and watch how the models behave.
@@ XI Heterogeneous Data, Missing Modalities, and Adversarial Attacks
Heterogeneous data and missing modalities pose significant challenges in multi-modal learning [2409.07825]. Factorized representations focus on shared structures when data is incomplete [1806.06176].
@@
where $\mathcal{M}_{\text{avail}}$ is the set of available modalities, $\mathbf{W}_{m}$ is the projection matrix, and $\mathbf{z}$ is the robust representation. Asymmetric reinforcement prioritizes stable modalities [2501.01240].
@@
in this equation, $\mathcal{L}_{m}$ is the loss for modality $m$, $w_{m}$ is its weight, and $\mathcal{L}_{\text{impute}}$ is the imputation loss. Adversarial attacks require robust defenses [1810.08437].
@@
Over the past few years, multi-modal learning has made impressive progress. We’ve seen strong performance in a wide range of domains, computer vision, natural language processing, healthcare, and more [2409.07825][1806.06176]. But still, let’s be honest: this field is far from “solved.” Several deep-rooted challenges continue to limit how far we can go. Some of them are technical; others are more about how the real world works. Either way, we need to face them head-on.
@@
One of the biggest headaches is dealing with heterogeneous data, that is, information coming from totally different sources, each with its own format and structure [2112.12792]. Imagine a dataset combining text, images, and audio. Each of these modalities has its own rhythm, its own kind of noise, its own “language.” And to expect a single model to handle them all smoothly? It is hard. The model must learn not just from each modality but also across them, and that’s where things start to break down.
@@
Another major challenge? Missing modalities. And trust me, this is not rare. A sensor might fail. Data gets corrupted. Or, in some cases, certain signals are deliberately excluded for privacy or cost reasons [2409.07825][2103.05677]. The result? A performance drop, sometimes a big one. So now, the question becomes: how do we build models that still function when part of the input goes missing? Some recent solutions offer promising ideas. For example, factorized representations help models focus on shared structures even when data is incomplete [1806.06176]. Others propose asymmetric reinforcement [2501.01240], guiding the model to rely more on stable signals. And then there are lightweight adaptation techniques [2310.03986], which offer practical, less resource-intensive fixes. These methods show one thing clearly: resilience is now a must.
@@
And resilience is not just about noise or gaps, it is also about adversarial attacks [1810.08437]. In high-stakes environments like medical AI or autonomous driving, even tiny, human-imperceptible changes to input data can lead to completely wrong predictions. That’s scary. So, defending against such attacks, detecting them, resisting them, recovering from them, has become a core concern [2411.14717]. Multi-modal models need to be not just smart, but safe and secure.
@@
Then there is the issue of modality imbalance, a subtle but serious problem. Sometimes one modality dominates the learning process because its signals are stronger or easier to model. This leads to the model “ignoring” weaker but still-important signals [2210.12562][2112.12792]. And that’s bad, because it reduces the diversity of information the model actually uses. So, what do we do? Some researchers suggest greedy modality selection [2210.12562], where the model picks the most useful signals step by step. Others turn to game-theoretic regularization [2411.07335], where learning is balanced like a strategic game between modalities. Both methods aim to reduce bias and boost fairness in training, something increasingly critical in today’s AI landscape.
@@
Looking across all this work, one thing is obvious: solving these core issues is key to building stronger, more reliable multi-modal systems [2409.07825][2304.04385]. It is not just about improving one metric here or there. It is about improving how we learn joint representations, how we deal with data loss, and how we protect against manipulation. And interestingly, many of the proposed techniques are cross-cutting. For example, factorized learning and asymmetric reinforcement are useful not just for missing data, but also for handling imbalance and improving robustness overall [1806.06176][2501.01240]. That kind of versatility is powerful.
@@
In summary, while the field of multi-modal learning has achieved notable progress, a range of persistent challenges continues to limit its broader adoption and effectiveness. Key technical obstacles, such as the integration of heterogeneous data, resilience to missing or corrupted modalities, and defense against adversarial interference, remain unsolved. Existing approaches, including factorized learning frameworks, robust training techniques, and lightweight adaptation methods, have demonstrated value across multiple problem domains rather than in isolation. This suggests that future research should focus on developing integrated solutions capable of addressing overlapping challenges within a unified framework. By understanding the interconnections between these problems, and by designing models that reflect such complexity, the field can move closer to producing truly robust, adaptive, and generalizable multi-modal systems suitable for deployment in real-world, imperfect environments [2409.07825][2304.04385].
@@ XII Future Directions in Multi-modal Research
As multi-modal learning continues to grow, clearer frameworks and practical taxonomies are increasingly critical [2408.14491, 2209.03430]. These tools help researchers manage the complexity of heterogeneous data and ambiguous integration rules [1705.09406, 2107.13782]. Structuring the field is as vital as advancing model design.
A key direction is the push for task-relative definitions of multi-modality, focusing on modalities relevant to specific tasks [2103.06304]. This approach enhances efficiency by prioritizing utility over exhaustive modeling, though challenges arise with unbalanced or misaligned modalities [2411.17040, 2406.05496]. Deep learning for co-learning and robust fusion strategies, such as (4) and (5), are gaining traction [2301.04856], enabling modalities to guide each other in challenging settings [2311.13165].
Despite progress, challenges persist. Combining diverse data types is inherently complex, with issues like unstructured, noisy, or missing data [2409.07825]. Large-scale labeled datasets, particularly for non-text modalities, remain scarce [2107.07502, 2405.19334]. These gaps hinder training flexible and reliable models, necessitating adaptive systems for imperfect conditions.
The field’s interdisciplinary nature, blending computer science, cognitive research, and beyond, will drive future growth. Flexible, real-world-ready solutions are essential, balancing theoretical advances with practical constraints like limited resources and incomplete labels. Ultimately, improving definitions, fusion strategies, and evaluation tools without overcomplicating systems will unlock multi-modal learning’s potential in AI, healthcare, and education [2408.14491].
@@ XIII The Evolving Landscape of Multi-modal Learning
The field of multi-modal learning has gone through noticeable changes in recent years. This shift has been largely driven by the rapid progress in areas like machine learning, computer vision, and natural language processing. According to studies such as [2408.14491][2103.06304][1705.09406][2209.03430][2301.04856], the field has grown into a dynamic research hotspot, with a wide range of tasks and challenges emerging along the way. A shared belief among many researchers is that combining different types of data, text, image, audio, or beyond, can offer a more complete and nuanced understanding of complex problems [2107.13782][2312.06037]. This idea has inspired the creation of new definitions, conceptual models, and frameworks to address the unique challenges that arise when working with multi-modal inputs [2103.06304][2309.12458].
@@
Across the literature, some recurring patterns begin to surface. A common one is the difficulty of handling multiple modalities: not just processing them in isolation, but figuring out how to connect them, align them, and model their interactions in a meaningful way [1705.09406][2209.03430]. At the same time, more and more researchers are turning to deep learning as a backbone for these efforts, applying it across domains such as healthcare, education, and autonomous systems [2301.04856][2304.04385]. In parallel, identifying open questions and mapping future directions has become a key feature of many recent surveys and reviews [2209.03430][2412.17759], a sign that the field is still very much evolving.
@@
Interestingly, there is no complete consensus on what multi-modality even means. Some researchers propose their own definitions, some build entirely new taxonomies, and others focus on different fusion strategies, early, late, mid-level, or even graph-based approaches. For instance, [2408.14491] provides a comprehensive literature review, while [2103.06304] introduces a redefined concept of modality interaction. These variations do not contradict each other so much as reflect the broad and multifaceted nature of the field.
@@
From a technical standpoint, the methods explored in these papers are also quite diverse. Some focus on theory and modeling, while others are more application-oriented. For example, [2110.14202] investigates meta-learning through a multi-task lens, and [2305.03125] proposes a strategy that alternates between increasing and decreasing inter-modal correlations to improve learning outcomes. All of this suggests that multi-modal learning is not a single discipline but a meeting point of many others, requiring interdisciplinary knowledge and flexibility in thinking.
@@
Still, despite all the enthusiasm, several core difficulties remain. There are calls for deeper theoretical work to solidify the foundations of this area [2209.03430][2309.12458]. Mixing and managing different types of data continues to be a technical bottleneck [1705.09406][2301.04856]. Moreover, training large multi-modal models often demands significant computational resources and access to large-scale, high-quality datasets [2301.04856][2412.17759]. Another concern that comes up again and again is the lack of transparency: many current models are still hard to interpret or explain in practice [2209.03430][2312.06037], which limits trust and usability in real-world settings.
@@
Overall, multi-modal learning has come a long way, and it still has a long way to go. Thanks to developments in machine learning and related fields, the tools we now have are more powerful and versatile than ever. But alongside this progress are real challenges, both technical and conceptual. What makes this area especially exciting is its potential: the promise of building smarter, more robust, and more interpretable models. With continued research and refinement, we can expect to see new breakthroughs and broader applications, especially in areas like computer vision, natural language processing, and human-computer interaction [2107.13782][2312.06037][2408.14491].